# Unlearning Inversion Attacks for Graph Neural Networks

### v1

@@ Abstract
Graph unlearning methods aim to efficiently remove the impact of sensitive data from trained GNNs without full retraining, assuming that deleted information cannot be recovered. In this work, we challenge this assumption by introducing the graph unlearning inversion attack: given only black-box access to an unlearned GNN and partial graph knowledge, can an adversary reconstruct the removed edges? We identify two key challenges: varying probability-similarity thresholds for unlearned versus retained edges, and the difficulty of locating unlearned edge endpoints, and address them with TrendAttack. First, we derive and exploit the confidence pitfall, a theoretical and empirical pattern showing that nodes adjacent to unlearned edges exhibit a large drop in model confidence. Second, we design an adaptive prediction mechanism that applies different similarity thresholds to unlearned and other membership edges. Our framework flexibly integrates existing membership inference techniques and extends them with trend features. Experiments on four real-world datasets demonstrate that TrendAttack significantly outperforms state-of-the-art GNN membership inference baselines, exposing a critical privacy vulnerability in current graph unlearning methods.
1 Introduction
@@
Graph-structured data is prevalent in numerous real-world applications, such as recommender systems wang2019neural ; he2020lightgcn ; zhang2024linear , social media platforms fan2019graph ; sankar2021graph ; liu2024score , and financial transaction networks dou2020enhancing ; liu2020alleviating ; wang2025enhance . Graph Neural Networks (GNNs) hamilton2017inductive ; velickovic2018graph ; wu2019simplifying have emerged as powerful tools for modelling such data, leveraging their ability to capture both node attributes and graph topology. The effectiveness of GNNs relies on the message-passing mechanism gilmer2017neural ; feng2022powerful , which iteratively propagates information between nodes and their neighbors. This enables GNNs to generate rich node representations, facilitating key tasks like node classification kipf2017semi ; lin2024trojan , link prediction zhang2018link ; wang2025bridging , and graph classification lee2018graph ; fu2021sagn .
@@
Despite their success, GNNs raise significant concerns about privacy risks due to the sensitive nature of graph data sajadmanesh2021locally ; dai2024comprehensive . Real-world datasets often contain private information, such as purchasing records in recommender systems zhang2021membership ; xin2023user or loan histories in financial networks pedarsani2011privacy ; wang2024data . During training, GNNs inherently encode such sensitive information into their model parameters. When these trained models are shared via model APIs, privacy breaches may occur. These risks have led to regulations like the GDPR mantelero2013eu , CCPA ccpa , and PIPEDA pipeda , which enforce the right to be forgotten, allowing users to request the removal of their personal data from systems and models. This demand has necessitated the development of methods to remove the influence of specific data points from trained GNNs, a process known as graph unlearning chien2022certified ; chen2022graph ; wu2023gif .
@@
A straightforward way of graph unlearning is to retrain the model from scratch on a cleaned training graph. However, this approach is computationally infeasible for large-scale graphs, such as purchase networks in popular e-commerce platforms wang2018billion ; jin2023amazon and social networks on social media armstrong2013linkbench ; ching2015one , which may involve billions of nodes and edges. To address this, a wide range of recent graph unlearning methods focus on efficient parameter manipulation techniques wu2023gif ; wu2023certified ; wu2024graphguard , which approximate the removal of data by adjusting model parameters based on their influence. These techniques are often regarded as privacy-preserving, assuming that the unlearned data cannot be reconstructed .
@@
Figure 1: Illustration of the unlearning inversion attack. Considering an online social network $\mathcal{G}_{\mathrm{orig}}$, where a user requests the deletion of sensitive friendship information, resulting in a cleaned graph $\mathcal{G}_{\mathrm{un}}$ and updated model parameters $\boldsymbol{\theta}_{\mathrm{un}}$. The GNN model may be shared with third-parties via black-box APIs. If an attacker, leveraging the model API and auxiliary information about $\mathcal{G}_{\mathrm{un}}$, can reconstruct the removed knowledge $\Delta\mathcal{G}$ through an unlearning inversion attack, sensitive relationships may be exposed, severely compromising user privacy.
@@
In this paper, we challenge the assumption that existing graph unlearning methods are robust to privacy attacks. Specifically, we investigate a novel and important problem of whether unlearned information can be recovered through an emerging family of membership inference attacks, namely unlearning inversion attacks hu2024learn (Figure 1). This attack is important because it reveals critical privacy vulnerabilities in graph unlearning methods and shows a concerning scenario where “forgotten” user information on the Web can be reconstructed through carefully designed attacks. Our central research question is: Can third-parties exploit model APIs of unlearned GNNs to recover sensitive information that was meant to be forgotten?
Although several pioneering studies have examined the privacy risks of unlearned ML models chen2021machine ; thudi2022necessity ; hu2024learn ; bertran2024reconstruction , conducting an unlearning inversion attack on GNNs presents unique technical challenges. Specifically, we argue that the following prior works may not fully resolve the proposed research question: (i) General-purpose unlearning inversion attacks hu2024learn ; bertran2024reconstruction often assume attacker’s access to model parameters bertran2024reconstruction or probability predictions hu2024learn before and after unlearning. However, this assumption may not hold in real-world settings, as the pre-unlearning model is highly sensitive and may not be accessible to third parties. (ii) Training a link prediction model directly on $\mathcal{G}_{\mathrm{un}}$ is an intuitive way to recover unlearned edges from the cleaned graph. However, due to API query limitations imposed by many model-serving platforms and online social networks (e.g., see query limitation policies of Twitter  [^1] https://developer.x.com/en/docs/x-api and TikTok  [^2] https://developers.tiktok.com/doc/research-api-faq), such direct link prediction techniques are less practical in real-world settings, where only partial access to $\mathcal{G}_{\mathrm{un}}$ is available. (iii) Membership inference attack (MIA) he2021node ; he2021stealing ; olatunji2021membership ; zhang2022inference methods designed for GNNs are another possible alternative. However, they are tailored to extract knowledge that is clearly “memorized” by the model, which is insufficient for unlearned GNNs that retain only minimal residual information of such sensitive knowledge.
@@
In this work, we study the link-level unlearning inversion attack problem for black-box unlearned GNN models, aiming to accurately recover unlearned links using black-box GNN outputs and partial knowledge of the unlearned graph $\mathcal{G}_{\mathrm{un}}$. We identify two key technical challenges for such attacks: (i) the output probabilities for two unlearned nodes may not exhibit sufficiently high similarity, which may require different prediction thresholds when inferring their membership compared to inferring other edges; (ii) it is highly non-trivial to determine whether a node is connected to an unlearned edge, since we only have access to the post-unlearning model, with no reference to pre-unlearning models. In response to these challenges, we propose a novel link-level unlearning inversion attack for black-box unlearned GNN models, namely TrendAttack. To address challenge (i), we design an adaptive prediction mechanism that applies different thresholds to infer two types of training graph edges: unlearned edges and other membership edges, enhancing TrendAttack’s flexibility to accommodate varying similarity levels. For challenge (ii), we identify a key phenomenon called the confidence pitfall, which enables the distinction between nodes connected to unlearned edges and others, using only black-box model outputs. This phenomenon describes how the model’s confidence in nodes near unlearned edges tends to decrease, and is supported by both empirical and theoretical evidence, as detailed in Section 5.1. By jointly incorporating the adaptive treatment of different edge types and the confidence trend pattern, TrendAttack achieves strong membership inference performance for both unlearned and other membership edges.
Our main contributions are: (i) We study a novel problem of graph unlearning inversion attack, pointing out the vulnerability of existing graph unlearning; (ii) We identify a simple yet effective pattern, the confidence pitfall, which distinguishes nodes connected to unlearned edges, supported by both empirical and theoretical evidence; (iii) We introduce a novel unlearning inversion attack TrendAttack that leverages confidence trend features in an adaptive membership inference framework, which can accurately identify unlearned edges and be flexibly integrated to existing GNN MIA methods; and (iv) Comprehensive evaluation on four real-world datasets shows that our method consistently outperforms state-of-the-art GNN MIA baselines in membership inference accuracy.
2 Related Works
@@
Graph Unlearning. Graph Unlearning enables the efficient removal of unwanted data’s influence from trained graph ML models chen2022graph ; said2023survey ; fan2025opengu . This removal process balances model utility, unlearning efficiency, and removal guarantees, following two major lines of research: retrain-based unlearning and approximate unlearning. Retrain-based unlearning partitions the original training graph into disjoint subgraphs, training independent submodels on them, enabling unlearning through retraining on a smaller subset of the data. Specifically, GraphEraser chen2022graph pioneered the first retraining-based unlearning framework for GNNs, utilizing balanced clustering methods for subgraph partitioning and ensembling submodels in prediction with trainable fusion weights to enhance model utility. Subsequently, many studies wang2023inductive ; li2023ultrare ; zhang2024graph ; li2025community have made significant contributions to improving the Pareto front of utility and efficiency in these methods, employing techniques such as data condensation li2025community and enhanced clustering li2023ultrare ; zhang2024graph . Approximate unlearning efficiently updates model parameters to remove unwanted data. Certified graph unlearning chien2022certified provides an important early exploration of approximate unlearning in SGC wu2019simplifying , with provable unlearning guarantees. GraphGuard wu2024graphguard introduces a comprehensive system to mitigate training data misuse in GNNs, featuring a significant gradient ascent unlearning method as one of its core components. GIF wu2023gif presents a novel influence function-based unlearning approach tailored to graph data, considering feature, node, and edge unlearning settings. Recent innovative works have further advanced the scalability pan2023unlearning ; li2024tcgu ; yi2025scalable ; yang2025erase ; zhang2025dynamic and model utility li2024towards ; zhang2025node of approximate unlearning methods. In this paper, we explore the privacy vulnerabilities of graph unlearning by proposing a novel membership inference attack tailored to unlearned GNN models, introducing a new defense frontier that graph unlearning should consider from a security perspective.
@@
Membership Inference Attack for GNNs. Membership Inference Attack (MIA) is a privacy attack targeting ML models, aiming to distinguish whether a specific data point belongs to the training set shokri2017membership ; hu2022membership . Recently, MIA has been extended to graph learning, where a pioneering work duddu2020quantifying explored the feasibility of membership inference in classical graph embedding models. Subsequently, interest has shifted towards attacking graph neural networks (GNNs), with several impactful and innovative studies revealing GNNs’ privacy vulnerabilities in node classification he2021node ; he2021stealing ; olatunji2021membership ; zhang2022inference and graph classification tasks wu2021adapting , covering cover node-level he2021node ; olatunji2021membership , link-level he2021stealing , and graph-level wu2021adapting ; zhang2022inference inference risks. Building on this, GroupAttack zhang2023demystifying presents a compelling advancement in link-stealing attacks he2021stealing on GNNs, theoretically demonstrating that different edge groups exhibit varying risk levels and require distinct attack thresholds, while a label-only attack has been proposed to target node-level privacy vulnerabilities conti2022label with a stricter setting. Another significant line of research involves graph model inversion attacks, which aim to reconstruct the graph structure using model gradients from white-box models zhang2021graphmi or approximated gradients from black-box models zhang2022model .
@@
Despite the impressive contributions of previous MIA studies in graph ML models, existing approaches overlook GNNs containing unlearned sensitive knowledge and do not focus on recovering such knowledge from unlearned GNN models. Additional related works are in Appendix A.
3 Preliminaries
@@
In this section, we present the notations used in this paper and give preliminaries on graph unlearning.
@@
Notations. In this paper, bold uppercase letters (e.g., $\mathbf{X}$) denote matrices, bold lowercase letters (e.g., $\mathbf{x}$) denote column vectors, and normal letters (e.g., $x$) indicate scalars. Let $\mathbf{e}_{u}\in\mathbb{R}^{d}$ be the column vector with the $u$-th element as $1$ and all others as 00. We use ||||| | to denote concatenating two vectors. We define the weighted inner product with a PSD matrix $\mathbf{H}$ as $\langle\mathbf{x},\mathbf{y}\rangle_{\mathbf{H}}:=\mathbf{x}^{\top}\mathbf{H}\mathbf{y}$. We use $\mathcal{A}\setminus\mathcal{B}:=\{x:x\in\mathcal{A},x\notin\mathcal{B}\}$ to denote the set difference between sets $\mathcal{A}$ and $\mathcal{B}$. Let $\mathcal{G}=(\mathcal{V},\mathcal{E})$ denote a graph, where $\mathcal{V}=\{v_{1},\cdots,v_{n}\}$ is the node set and $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$ is the edge set. The node feature of node $v_{i}\in\mathcal{V}$ is denoted by $\mathbf{x}_{i}\in\mathbb{R}^{d}$, and the feature matrix for all nodes is denoted by $\mathbf{X}=[\mathbf{x}_{1},\cdots,\mathbf{x}_{n}]^{\top}\in\mathbb{R}^{n\times d}$. The adjacency matrix $\mathbf{A}\in\{0,1\}^{n\times n}$ encodes the edge set, where $A_{i,j}=1$ if $(v_{i},v_{j})\in\mathcal{E}$, and $\mathbf{A}_{i,j}=0$ otherwise. Specifically, $\mathbf{D}\in\mathbb{R}^{n\times n}$ is the degree matrix, where the diagonal elements $D_{i,i}=\sum_{j=1}^{n}A_{i,j}$. We use $\mathcal{N}(v)$ to denote the neighborhood of node $v_{i}\in\mathcal{V}$, and use $\widehat{\mathcal{N}}(v)$ to denote a subset of $\mathcal{N}(v)$. Specifically, $\mathcal{N}^{(k)}(v_{i})$ represents all nodes in $v_{i}$’s $k$-hop neighborhood.
@@
Semi-supervised Node Classification. We focus on a semi-supervised node classification task in a transductive setting, which is common in real-world applications kipf2017semi ; wu2020comprehensive . In this setting, the training graph $\mathcal{G}$ includes a small subset of labeled nodes $\mathcal{V}_{L}=\{u_{1},\cdots,u_{|\mathcal{V}_{L}|}\}\subseteq\mathcal{V}$, where each node is annotated with a label $y\in\mathcal{Y}$. The remaining nodes are unlabeled and belong to the subset $\mathcal{V}_{U}$, where $\mathcal{V}_{U}\cap\mathcal{V}_{L}=\emptyset$. The test set $\mathcal{V}_{T}$ is a subset of the unlabeled nodes, represented as $\mathcal{V}_{T}\subseteq\mathcal{V}_{U}$. We denote the GNN output for a specific target node $v\in\mathcal{V}$ as $f_{\mathcal{G}}(v;\boldsymbol{\theta})$, where $\mathcal{G}$ is the graph used for neighbor aggregation and $\theta$ is the model parameters.
@@
Graph Unlearning. Graph unlearning aims to remove the impact of some undesirable training data from trained GNN models under limited computational overhead chen2022graph ; wu2023gif . Specifically, consider the original training graph $\mathcal{G}_{\mathrm{orig}}:=(\mathcal{V}_{\mathrm{orig}},\mathcal{E}_{\mathrm{orig}})$ including both desirable data and undesirable data. Normally, the parameters $\boldsymbol{\theta}_{\mathrm{orig}}$ of the GNN model trained on the original graph is given as:
$\small\boldsymbol{\theta}_{\mathrm{orig}}:=\mathop{\arg\min}_{\boldsymbol{\theta}}{\sum}_{v\in\mathcal{V}_{L}}\mathcal{L}(f_{\mathcal{G}_{\mathrm{orig}}}(v;\boldsymbol{\theta}),y_{v}),$ (1)
@@
where $\mathcal{V}_{L}\subseteq\mathcal{V}_{\mathrm{orig}}$ is the set of labeled nodes, and $\mathcal{L}$ is an loss function (e.g., cross-entropy zhang2018generalized ).
@@
Let the undesirable knowledge be a subgraph $\Delta\mathcal{G}:=(\Delta\mathcal{V},\Delta\mathcal{E})$ of the original graph, where $\Delta\mathcal{V}\subseteq\mathcal{V}_{\mathrm{orig}}$ and $\Delta\mathcal{E}\subseteq\mathcal{E}_{\mathrm{orig}}$. The unlearned graph is defined as $\mathcal{G}_{\mathrm{un}}:=(\mathcal{V}\setminus\Delta\mathcal{V},\mathcal{E}\setminus\Delta\mathcal{E})$, which excludes the undesirable knowledge. The goal of graph unlearning is to obtain parameters $\boldsymbol{\theta}_{\mathrm{un}}$ using an efficient algorithm Unlearn (e.g., gradient ascent wu2024graphguard ; zhang2025catastrophic , or influence function computation wu2023gif ; wu2023certified ), such that $\boldsymbol{\theta}_{\mathrm{un}}$ closely approximates the retrained parameters $\boldsymbol{\theta}_{\mathrm{re}}$ from the cleaned graph $\mathcal{G}_{\mathrm{un}}$, while being significantly more efficient than retraining from scratch. Formally, the unlearning process is defined as:
@@
$\small\boldsymbol{\theta}_{\mathrm{un}}:=~{}\textsc{Unlearn}(f,\boldsymbol{\theta}_{\mathrm{orig}},\mathcal{G}_{\mathrm{orig}},\Delta\mathcal{G})\approx~{}\mathop{\arg\min}_{\boldsymbol{\theta}}{\sum}_{v\in\mathcal{V}_{L}/\Delta\mathcal{V}}\mathcal{L}(f_{\mathcal{G}_{\mathrm{orig}}/\Delta\mathcal{G}}(v;\boldsymbol{\theta}),y_{v}),$ (2)
where the right optimization problem denotes retrain from scratch on the unlearned graph.

### v2

@@ Abstract.
Graph unlearning methods aim to efficiently remove the impact of sensitive data from trained GNNs without full retraining, assuming that deleted information cannot be recovered. In this work, we challenge this assumption by introducing the graph unlearning inversion attack: given only black-box access to an unlearned GNN and partial graph knowledge, can an adversary reconstruct the removed edges? We identify two key challenges: varying probability-similarity thresholds for unlearned versus retained edges, and the difficulty of locating unlearned edge endpoints, and address them with TrendAttack. First, we derive and exploit the confidence pitfall, a theoretical and empirical pattern showing that nodes adjacent to unlearned edges exhibit a large drop in model confidence. Second, we design an adaptive prediction mechanism that applies different similarity thresholds to unlearned and other membership edges. Our framework flexibly integrates existing membership inference techniques and extends them with trend features. Experiments on four real-world datasets demonstrate that TrendAttack significantly outperforms state-of-the-art GNN membership inference baselines, exposing a critical privacy vulnerability in current graph unlearning methods. The source code is available on GitHub [^1] https://github.com/QwQ2000/WSDM26-Graph-Unlearning-Inversion.
Graph Unlearning, Privacy Attack, Graph Neural Networks
copyright: nonecopyright: acmlicensedjournalyear: 2026copyright: ccconference: Proceedings of the Nineteenth ACM International Conference on Web Search and Data Mining; February 22–26, 2026; Boise, ID, USAbooktitle: Proceedings of the Nineteenth ACM International Conference on Web Search and Data Mining (WSDM ’26), February 22–26, 2026, Boise, ID, USAdoi: 10.1145/3773966.3777929isbn: 979-8-4007-2292-9/2026/02ccs: Computing methodologies Machine learningccs: Security and privacy Privacy protections
1. Introduction
@@
Graph-structured data is prevalent in numerous real-world applications, such as recommender systems (he2020lightgcn; zhang2024linear; liu2026continuous), social media platforms (fan2019graph; sankar2021graph; liu2024score), financial transaction networks (dou2020enhancing; liu2020alleviating; wang2025enhance), and computational biology (jing2021fast; fan2025computational; xu2025dualequinet). Graph Neural Networks (GNNs) (hamilton2017inductive; velickovic2018graph; wu2019simplifying) have emerged as powerful tools for modelling such data, leveraging their ability to capture both node attributes and graph topology. The effectiveness of GNNs relies on the message-passing mechanism (gilmer2017neural; feng2022powerful), which iteratively propagates information between nodes and their neighbors. This enables GNNs to generate rich node representations, facilitating key tasks like node classification (kipf2017semi; xue2023lazygnn; lin2024trojan), link prediction (zhang2018link; yang2024conversational; wang2025bridging), and graph classification (lee2018graph; fu2021sagn; al2025graph).
@@
Despite their success, GNNs raise significant concerns about privacy risks due to the sensitive nature of graph data (sajadmanesh2021locally; dai2024comprehensive). Real-world datasets often contain private information, such as purchasing records in recommender systems (zhang2021membership; xin2023user) or loan histories in financial networks (pedarsani2011privacy; wang2024data). During training, GNNs inherently encode such sensitive information into their model parameters. When these trained models are shared via model APIs, privacy breaches may occur. These risks have led to regulations like the GDPR (mantelero2013eu), CCPA (ccpa), and PIPEDA (pipeda), which enforce the right to be forgotten, allowing users to request the removal of their personal data from systems and models. This demand has necessitated the development of methods to remove the influence of specific data points from trained GNNs, a process known as graph unlearning (chien2022certified; chen2022graph; wu2023gif).
@@
A straightforward way of graph unlearning is to retrain the model from scratch on a cleaned training graph. However, this approach is computationally infeasible for large-scale graphs, such as purchase networks in popular e-commerce platforms (wang2018billion; jin2023amazon) and social networks on social media (armstrong2013linkbench; ching2015one), which may involve billions of nodes and edges. To address this, a wide range of recent graph unlearning methods focus on efficient parameter manipulation techniques (wu2023gif; wu2023certified; wu2024graphguard), which approximate the removal of data by adjusting model parameters based on their influence. After unlearning, membership inference attacks for GNNs (olatunji2021membership; he2021stealing; zhang2023demystifying) could become unable to determine whether the unlearned edges were part of the training set, as the GNN no longer memorizes them. Hence, these techniques are often regarded as privacy-preserving, assuming that the unlearned data cannot be reconstructed.
@@
Figure 1. Illustration of the unlearning inversion attack. Considering an online social network $\mathcal{G}_{\mathrm{orig}}$, where a user requests the deletion of sensitive friendship information, resulting in a cleaned graph $\mathcal{G}_{\mathrm{un}}$ and updated model parameters $\bm{\theta}_{\mathrm{un}}$. The GNN model may be shared with third-parties via black-box APIs. If an attacker, leveraging the model API and auxiliary information about $\mathcal{G}_{\mathrm{un}}$, can reconstruct the removed knowledge $\Delta\mathcal{G}$ through an unlearning inversion attack, sensitive relationships may be exposed, severely compromising user privacy.
@@
In this paper, we challenge the assumption that existing graph unlearning methods are robust to privacy attacks. Specifically, we study a novel and important problem of whether unlearned information can be recovered through unlearning inversion attacks (hu2024learn) as illustrated in Figure 1. This attack is important because it reveals critical privacy vulnerabilities in graph unlearning methods and shows a concerning scenario where “forgotten” user information on the Web can be reconstructed through carefully designed attacks. Our central research question is:
Can third parties exploit model APIs of unlearned GNNs to recover sensitive information that was meant to be forgotten?
Although several pioneering studies have examined the privacy risks of unlearned machine learning (ML) models (chen2021machine; thudi2022necessity; hu2024learn; bertran2024reconstruction), unlearning inversion attacks on GNNs still remain largely unexplored. Existing general-purpose unlearning inversion attacks (hu2024learn; bertran2024reconstruction) typically adopt a “two-model” setting, where the attacker is assumed to have access to model parameters (bertran2024reconstruction) or output probabilities (hu2024learn) both before and after unlearning. This assumption is often unrealistic in practice, as pre-unlearning models may contain sensitive information and are usually inaccessible to third parties, necessitating a more practical “one-model” setting where the attacker only has access to the model after unlearning. Moreover, these methods are designed for i.i.d. data (e.g., images) and fail to capture the non-i.i.d. nature of graphs, including structural dependencies and the propagation of unlearning effects across the network. It is also worth noting that other privacy attacks on GNNs, such as membership inference attacks (MIAs) (he2021node; he2021stealing; olatunji2021membership; zhang2022inference), may also be less effective in this context, as they aim to extract knowledge explicitly memorized by the model, while unlearned GNNs have already forgotten most of this sensitive information.
Therefore, in this work, we study a novel problem of link-level unlearning inversion attack for black-box unlearned GNN models, with the goal of accurately recovering unlearned links using only black-box GNN outputs and partial knowledge of the unlearned graph $\mathcal{G}_{\mathrm{un}}$. A straightforward solution is to train a link classifier on the output probabilities of the unlearned GNN model $f(\cdot;\bm{\theta}_{\mathrm{un}})$ to decide whether an edge belongs to the training data (regardless of whether the edge was unlearned). However, this naive approach faces two major technical challenges: (i) The output probabilities of two nodes connected by an unlearned edge often do not show a strong similarity signal, making these unlearned edges hard to separate from ordinary non-existent edges. (ii) Even if we could measure such similarity, it is fundamentally difficult to identify which nodes are associated with unlearned edges, because we only observe the post-unlearning model and have no access to the pre-unlearning model as a reference.
@@
In response to these challenges, we propose a novel link-level unlearning inversion attack for black-box unlearned GNN models, namely TrendAttack. To address challenge (i), we design an adaptive prediction mechanism that applies different thresholds to infer two types of training graph edges: unlearned edges and other membership edges, enhancing TrendAttack’s flexibility to accommodate varying similarity levels. For challenge (ii), we identify a key phenomenon called the confidence pitfall, which enables the distinction between nodes connected to unlearned edges and others, using only black-box model outputs. This phenomenon describes how the model’s confidence in nodes near unlearned edges tends to decrease, and is supported by both empirical and theoretical evidence, as detailed in Section 5.1. By jointly incorporating the adaptive treatment of different edge types and the confidence trend pattern, TrendAttack achieves strong membership inference performance for both unlearned and other membership edges. Our main contributions are summarized as follows:
  • We formulate a novel problem of graph unlearning inversion attacks (Section 4), highlighting the vulnerability of existing graph unlearning methods. This work is among the first to study GNN unlearning vulnerabilities via privacy attacks.
  • We identify a simple yet effective pattern, the confidence pitfall (Section 5.1), which distinguishes nodes connected to unlearned edges, supported by both empirical and theoretical evidence.
  • We introduce a novel unlearning inversion attack, namely TrendAttack (Section 5.2), which leverages confidence pitfall to accurately identify unlearned edges from black-box GNN outputs.
  • Comprehensive evaluation (Section 6) on four real-world datasets shows that our method consistently outperforms state-of-the-art GNN privacy attacks on recovering unlearned edges.
2. Related Works
@@
Graph Unlearning. Graph Unlearning enables the efficient removal of unwanted data’s influence from trained graph ML models (chen2022graph; said2023survey; fan2025opengu). This removal process balances model utility, unlearning efficiency, and removal guarantees, following two major lines of research: retrain-based unlearning and approximate unlearning. Retrain-based unlearning partitions the original training graph into disjoint subgraphs, training independent submodels on them, enabling unlearning through retraining on a smaller subset of the data. Specifically, GraphEraser (chen2022graph) pioneered the first retraining-based unlearning framework for GNNs, utilizing balanced clustering methods for subgraph partitioning and ensembling submodels in prediction with trainable fusion weights to enhance model utility. Subsequently, many studies (wang2023inductive; li2023ultrare; zhang2024graph; li2025community) have made significant contributions to improving the Pareto front of utility and efficiency in these methods, employing techniques such as data condensation (li2025community) and enhanced clustering (li2023ultrare; zhang2024graph). Approximate unlearning efficiently updates model parameters to remove unwanted data. Certified graph unlearning (chien2022certified) provides an important early exploration of approximate unlearning in SGC (wu2019simplifying), with provable unlearning guarantees. GraphGuard (wu2024graphguard) introduces a comprehensive system to mitigate training data misuse in GNNs, featuring a significant gradient ascent unlearning method as one of its core components. GIF (wu2023gif) presents a novel influence function-based unlearning approach tailored to graph data, considering feature, node, and edge unlearning settings. Recent innovative works have further advanced the scalability (pan2023unlearning; li2024tcgu; yi2025scalable; yang2025erase; zhang2025dynamic) and model utility (li2024towards; zhang2025node) of approximate unlearning methods. In this paper, we explore the privacy vulnerabilities of graph unlearning by proposing a novel membership inference attack tailored to unlearned GNN models, introducing a new defense frontier that graph unlearning should consider from a security perspective.
@@
Membership Inference Attack for GNNs. Membership Inference Attack (MIA) is a privacy attack targeting ML models, aiming to distinguish whether a specific data point belongs to the training set (shokri2017membership; hu2022membership). Recently, MIA has been extended to graph learning, where a pioneering work (duddu2020quantifying) explored the feasibility of membership inference in classical graph embedding models. Subsequently, interest has shifted towards attacking graph neural networks (GNNs), with several impactful and innovative studies revealing GNNs’ privacy vulnerabilities in node classification (he2021node; he2021stealing; olatunji2021membership; zhang2022inference) and graph classification tasks (wu2021adapting), covering cover node-level (he2021node; olatunji2021membership), link-level (he2021stealing), and graph-level (wu2021adapting; zhang2022inference) inference risks. Building on this, GroupAttack (zhang2023demystifying) presents a compelling advancement in link-stealing attacks (he2021stealing) on GNNs, theoretically demonstrating that different edge groups exhibit varying risk levels and require distinct attack thresholds, while a label-only attack has been proposed to target node-level privacy vulnerabilities (conti2022label) with a stricter setting. Another significant line of research involves graph model inversion attacks, which aim to reconstruct the graph structure using model gradients from white-box models (zhang2021graphmi) or approximated gradients from black-box models (zhang2022model).
@@
Despite the impressive contributions of previous MIA studies in graph ML models, existing approaches overlook GNNs containing unlearned sensitive knowledge and do not focus on recovering such knowledge from unlearned GNN models. Additional related works are in Supplementary B.
3. Preliminaries
@@
In this section, we present the notations used in this paper and give preliminaries on node classification and graph unlearning.
@@
Notations. In this paper, bold uppercase letters (e.g., $\mathbf{X}$) denote matrices, bold lowercase letters (e.g., $\mathbf{x}$) denote column vectors, and normal letters (e.g., $x$) indicate scalars. We use $\mathcal{A}\setminus\mathcal{B}:=\{x:x\in\mathcal{A},x\notin\mathcal{B}\}$ to denote the set difference between sets $\mathcal{A}$ and $\mathcal{B}$. Let $\mathcal{G}=(\mathcal{V},\mathcal{E})$ denote a graph, where $\mathcal{V}=\{v_{1},\cdots,v_{n}\}$ is the node set and $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$ is the edge set. The node feature of node $v_{i}\in\mathcal{V}$ is denoted by $\mathbf{x}_{i}\in\mathbb{R}^{d}$, and the feature matrix for all nodes is denoted by $\mathbf{X}=[\mathbf{x}_{1},\cdots,\mathbf{x}_{n}]^{\top}\in\mathbb{R}^{n\times d}$. The adjacency matrix $\mathbf{A}\in\{0,1\}^{n\times n}$ encodes the edge set, where $A_{i,j}=1$ if $(v_{i},v_{j})\in\mathcal{E}$, and $\mathbf{A}_{i,j}=0$ otherwise. Specifically, $\mathbf{D}\in\mathbb{R}^{n\times n}$ is the degree matrix, where the diagonal elements $D_{i,i}=\sum_{j=1}^{n}A_{i,j}$. We use $\mathcal{N}(v)$ to denote the neighborhood of node $v_{i}\in\mathcal{V}$, and use $\widehat{\mathcal{N}}(v)$ to denote a subset of $\mathcal{N}(v)$. Specifically, $\mathcal{N}^{(k)}(v_{i})$ represents all nodes in $v_{i}$’s $k$-hop neighborhood. A full list of notations is in Supplementary A.
Semi-supervised Node Classification. We focus on a semi-superv
@@
-ised node classification task in a transductive setting, which is common in real-world applications (kipf2017semi; wu2020comprehensive). In this setting, the training graph $\mathcal{G}$ includes a small subset of labeled nodes $\mathcal{V}_{L}=\{u_{1},\cdots,u_{|\mathcal{V}_{L}|}\}\subseteq\mathcal{V}$, where each node is annotated with a label $y\in\mathcal{Y}$. The remaining nodes are unlabeled and belong to the subset $\mathcal{V}_{U}$, where $\mathcal{V}_{U}\cap\mathcal{V}_{L}=\emptyset$. The test set $\mathcal{V}_{T}$ is a subset of the unlabeled nodes, represented as $\mathcal{V}_{T}\subseteq\mathcal{V}_{U}$. We denote the GNN output for a specific target node $v\in\mathcal{V}$ as $f_{\mathcal{G}}(v;\bm{\theta})$, where $\mathcal{G}$ is the graph used for neighbor aggregation and $\theta$ is the model parameters.
@@
Graph Unlearning. Graph unlearning aims to remove the impact of some undesirable training data from trained GNN models under limited computational overhead (chen2022graph; wu2023gif). Specifically, consider the original training graph $\mathcal{G}_{\mathrm{orig}}:=(\mathcal{V}_{\mathrm{orig}},\mathcal{E}_{\mathrm{orig}})$ including both desirable data and undesirable data. Normally, the parameters $\bm{\theta}_{\mathrm{orig}}$ of the GNN model trained on the original graph is given as:
(1) $\bm{\theta}_{\mathrm{orig}}:=\mathop{\arg\min}_{\bm{\theta}}\sum_{v\in\mathcal{V}_{L}}\mathcal{L}(f_{\mathcal{G}_{\mathrm{orig}}}(v;\bm{\theta}),y_{v}),$
@@
where $\mathcal{V}_{L}\subseteq\mathcal{V}_{\mathrm{orig}}$ is the set of labeled nodes, and $\mathcal{L}$ is a loss function for node classification (e.g., cross-entropy (zhang2018generalized)).
@@
Let the undesirable knowledge be a subgraph $\Delta\mathcal{G}:=(\Delta\mathcal{V},\Delta\mathcal{E})$ of the original graph, where $\Delta\mathcal{V}\subseteq\mathcal{V}_{\mathrm{orig}}$ and $\Delta\mathcal{E}\subseteq\mathcal{E}_{\mathrm{orig}}$. The unlearned graph is defined as $\mathcal{G}_{\mathrm{un}}:=(\mathcal{V}\setminus\Delta\mathcal{V},\mathcal{E}\setminus\Delta\mathcal{E})$, which excludes the undesirable knowledge. The goal of graph unlearning is to obtain parameters $\bm{\theta}_{\mathrm{un}}$ using an efficient algorithm Unlearn (e.g., gradient ascent (wu2024graphguard; zhang2025catastrophic), or influence function computation (wu2023gif; wu2023certified)), such that $\bm{\theta}_{\mathrm{un}}$ closely approximates the retrained parameters $\bm{\theta}_{\mathrm{re}}$ from the cleaned graph $\mathcal{G}_{\mathrm{un}}$, while being significantly more efficient than retraining from scratch, i.e.,
$\bm{\theta}_{\mathrm{un}}:=$ $~\textsc{Unlearn}(f,\bm{\theta}_{\mathrm{orig}},\mathcal{G}_{\mathrm{orig}},\Delta\mathcal{G})$
@@
(2) $\approx$ $~\mathop{\arg\min}_{\bm{\theta}}\sum_{v\in\mathcal{V}_{L}\setminus\Delta\mathcal{V}}\mathcal{L}(f_{\mathcal{G}_{\mathrm{orig}}\setminus\Delta\mathcal{G}}(v;\bm{\theta}),y_{v}),$